MONT 104Q – Mathematical Journeys First Writing Assignment

The first of the larger paper assignments this semester will be devoted to the Odyssey. It is also designed to establish a baseline for your current level of mastery of an important form of writing sometimes called the argumentative essay. Note: “argumentative” here does not mean that you are trying to pick a fight through your writing(!) It means that you are making a a particular response to a question – a thesis – and then presenting appropriate evidence and reasoning to support your claim(s). Your essay should show that you have done some careful reading, and some serious thinking, about the subject. You may, of course, disagree with ideas presented in class if you want. The point is for you to develop your own thinking, and then clearly express it in writing.

Your essay should be 2-3 pages in length (double-spaced, single-sided). Please submit as a MS Word .doc or .docx file to jlittle@holycross.edu by 5:00pm on Friday, September 18.

The topic and question(s) you should address is this:

It is often the case that epic journeys in literature change the main character(s) in some way, teach them important lessons, or even turn them into radically different people than they were at the start. But is any of that true of Odysseus? Is he a different person at the end of this story than he was when he left for Troy? Have the Trojan War and all the experiences he has on his long journey back to Ithaca changed him in any appreciable way? Have some of them had that sort of effect, but not others? If that is true, which experiences have changed him and how? If, on the other hand, you think Odysseus regains his home and wife but is completely unchanged by his journey, what is the point of the whole story?

Guidelines for this assignment

Follow these specifications:

1. It is OK to discuss preliminary ideas for what you want to say with your classmates if you want. I will also be happy to talk over things and/or read a preliminary draft to give feedback if you like.

2. In the paper itself, develop, and then clearly state your thesis in your opening paragraph. The thesis should be a good answer to the main question: Is Odysseus a different person at the end of this story than he was when he left for Troy? Have the Trojan War and all the experiences he has on his long journey back to Ithaca changed him in any appreciable way? The thesis should be clearly stated and well focused. The reader (me) should not need to guess what your point is.

3. In the following paragraphs, using sound reasoning, present your evidence for making the claim in your thesis. Do not contradict what you said earlier. Do not ignore evidence that contradicts your ideas. If, on further reflection, you find that you do not agree with what you said before, revise the thesis(!)

4. In this sort of writing, a “move” that is often very effective is to temporarily entertain an opposing conclusion as a possibility, then explain the evidence that leads you to reject that conclusion. Try it!

5. Your evidence should consist of passages from the Lombardo translation of the Odyssey, and interpretations of those passages that come from your own thinking about them. Your evidence should be as specific, detailed, correct, and as varied as possible (focus on small passages and mention the little things that support your point).

6. Do not consult any other sources beyond the Lombardo translation of the Odyssey – this means no other translations, other books, no web sites, etc. When you use direct quotations from Lombardo’s translation, at the end place the book and line number(s) to identify the source. Here’s an example of what I’m looking for:

Odysseus tends to judge others by how clever and self-reliant they are. For instance, when he tells the Phaeacian Alcinous about his journey up to that point, he calls the Cyclopes “lawless savages who leave everything up to the gods” (Book 9, 105-106).

7. The paper as a whole should be well-organized and “flow” from one point to the next. Do not repeat pieces of evidence unless there is a very good reason for doing so. Do not get ahead of yourself. Say one thing at a time. If you decide you want to add something, do not just tack it on somewhere. Instead, figure out where it fits in the overall scheme of your argument and put it where it belongs.

8. Your sentences should be clear, readable, and not overly wordy. Aim for a reasonably formal but lively style. Careless or colloquial ways of saying things that are acceptable in casual conversation are not appropriate in this kind of writing. Also, please do not use a thesaurus to find impressive-sounding words when simpler alternatives are available. An important lesson to learn for writing in college is that thesauri do not capture the nuances in meanings that underlie good writing. If you must use a thesaurus, also look up any words you select in a dictionary to make sure they mean exactly what you think they do!

9. Errors in mechanics and usage should not annoy the reader. (As a warning, I tend to be easily annoyed by this sort of thing!) Try as much as possible to avoid mistakes in grammar, spelling, punctuation, and word choice.